OpenStudy (anonymous):

Simplify. n^6 · n^5 · n^4 ÷ n^3 · n^2 ÷ n isn't the answer n^9?

OpenStudy (anonymous):

No I don't think so..

OpenStudy (anonymous):

can you help me to figure it out?

OpenStudy (anonymous):

Solve one by one: \[\large n^6 .n^5 = ??\]

OpenStudy (anonymous):

n^11?

OpenStudy (anonymous):

Yes.. Now: \[\large \frac{n^4}{ n^3} = ??\]

OpenStudy (anonymous):

Just subtract the powers..

OpenStudy (anonymous):

n^1?

OpenStudy (anonymous):

Yes.. \[\large \frac{n^2}{ n} = ??\]

OpenStudy (anonymous):

n^1?

OpenStudy (anonymous):

So combine all the results now what did you get on combining??

OpenStudy (anonymous):

Combine all the three results you got...

OpenStudy (anonymous):

\[\large n^{11} \times n^1 \times n^1 = ???\]

OpenStudy (jiteshmeghwal9):

Use the formula BODMAS B=BRACKET O=OF D=DIVISION M=MULTIPLICATION A=ADDITION S=SUBTRACTION

OpenStudy (anonymous):

n^13?

OpenStudy (anonymous):

Yes well done..

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

Welcome dear..